OpenStudy (anonymous):

What is the solution to the system below? 2x+y=3 4x+2y=6 a(2, 2) b(2, -2) cno solution dinfinite solutions on the line

OpenStudy (anonymous):

First you would multiply the first equation by -2

OpenStudy (anonymous):

-2(2x + y)=-2(3)

OpenStudy (anonymous):

-4x + -2y=-6

OpenStudy (anonymous):

-4x + -2y=-6 +4x + 2y=6

OpenStudy (anonymous):

=0=0

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

thank u your the best :)